OpenStudy (surryyy):

Given that a truck can drive 25 miles with 4 gallons of gasoline, how many gallons does it need to drive 125 miles? 5 gallons 10 gallons 20 gallons 25 gallons

OpenStudy (steve816):

20 gallons

OpenStudy (sohailiftikhar):

as given 25 miles = 4 galones now 5 * 25 = 5 * 4 125 miles = 20 galones

OpenStudy (steve816):

galones? lol xD

OpenStudy (sohailiftikhar):

heheeh